OpenStudy (anonymous):

solve the substitution method 4x+3y=-17 x=22-6y

OpenStudy (anonymous):

4x+3y=-17x 21x+3y=0 -17x=22-6y -17x=22-2(-21x) -17x-42x=22 -59x=22 x=-22/59

OpenStudy (anonymous):

what is it in ordered pair?

OpenStudy (anonymous):

(-22/59, 154/59)

OpenStudy (anonymous):

did u got it